OpenStudy (anonymous):

Solve the following system. y = x + 3 3x + y = 19

OpenStudy (anonymous):

6.5

OpenStudy (goformit100):

You can use Matrices as Before :)

OpenStudy (uri):

Arrange the first eq first

OpenStudy (goformit100):

what then ?

OpenStudy (goformit100):

@uri reply ?

OpenStudy (uri):

y = x + 3 x+y=-3

OpenStudy (goformit100):

ok done then ?

OpenStudy (uri):

x+y=-3 3x+y=19(mutiply by -1) x+y=-3 3x-y=19 ======= 4x=16 x=16/4 x=4 We have value of x now find y

OpenStudy (anonymous):

is it -7 @uri

OpenStudy (anonymous):

or 7 @uri

OpenStudy (carson889):

Another way you could have done it is to solve both equations for a single variable (for example y): Thus: y = x + 3 and y = 19 - 3x Next set the y variables equal to each other by placing one on one side of the equal sign and the other on the other as such: x + 3 = 19 - 3x Next you want to group like terms, so group the x's together on one side and the numbers on the other: x + 3 = 19 - 3x x + 3x + 3 = 19 -3x + 3x x + 3x + 3 - 3 = 19 - 3 x + 3x = 19 - 3 Simplify: 4x = 16 x =16/4 = 4 To solve for why plug the value we just calculated for x into one of the original equations: y = x + 3 --> y = 4 + 3 = 7 To double check,use the other equation: 3x + y = 19 3*4 + y = 19 12 + y = 19 Re-arrange y = 19-12 y = 7 Bam!

OpenStudy (uri):

I did it right,Did not I?

OpenStudy (uri):

Put the value of x in original eq no.2 3x + y =19 3(4) +y =19 12+y=19 y=19-12 y=7

OpenStudy (carson889):

@uri Yes your analysis is completely right

OpenStudy (uri):

:3